CAT Chat Minutes

January 29, 2007

Operations Update, Mark Beno

The summer (2007-2) run is posted on the website. As a result of the continuing resolution, there will be a total of 4 weeks extra shutdown. This estimate is determined by eliminating one week of beamtime in April, one week at the end of May, one week in August, and one week at the end of September. If the budget is passed, these weeks may not be eliminated at all.

ESH Update, Tom Barkalow

An incident where an individual was left in a free-electron laser station was documented in a recent lessons learned report from Jefferson Lab. This individual heard the alarm, and was able to inform the person running the station. Please take a look at your sweep procedures and your station layout to ensure that procedures are in order and that a similar situation may be avoided.

As a reminder, machine shops will be surveyed for lead in March. The last survey was conducted late last year. There were two no lead areas that contained lead. Also, all machine shops must be cleaned after each use.

There were two incidents in December that caught M. Gibson's attention. One was a splash of lead nitrate in an eye, and a finger laceration. Particular areas mandate eye protection. Clarifications of these areas are in process now. Signage and glasses will be provided in areas not yet posted. A survey questionnaire will be sent out to identify what is being done to protect an individual from pinch point areas.

The DOE Worker Safety Rule, 10CFR, was issued last year and will go into effect in a couple of months. This rule will impact our user operations. This rule mandates certain standards be followed. Fines are applicable if there is non-compliance. A talk will be given to explain this rule in more detail.

APS Lattice Information Update, Michael Borland

The standard APS lattice has changed this run. The schedule page will be updated to include more information about lattices for present and past runs. Based on experience from the last run, we had two operational readiness periods during start-up. This program was put into place in an attempt to detect operational problems prior to the start of beamtime. Several beamlines have reported seeing an x-y tilt on the beam. Progress has been made on this issue and plans have been implemented to get them corrected. For further details on lattice changes, operational readiness, and beam tilt issues, please see the attached document.

User Input for Future ANL ES&H and Infrastructure Projects, John MacLean

Argonne maintains a list of ESH & Infrastructure projects required at the lab (http://www.tis.anl.gov/db/doccenter/full_document/DDD/202.pdf). We need to make sure any APS requirements get on the list.

Recent APS projects include:

  • Increase in chiller capacity
  • Assistance with LOM 437 construction
  • VESDA upgrade

APS listed projects include:

  • Further increase in chiller capacity
  • LOM435 parking lot expansion
  • LOM436 parking lot expansion
  • Revision of APS site signage
  • Resurface Rock and Kearny Roads
  • Johnson Controls network replacement
If you have any suggestions please email John MacLean at jfm@aps.anl.gov.

Operations Info from Shutdown Work: ID Work, X-ray BPMs, and Switchgear Maintenance, John Quintana

Switchgears need to shutdown once every three years for 12 hours for maintenance and inspections. This normally takes place over a weekend. Three switchgears are scheduled for maintenance during the April/May shutdown period. The 12-hour shutdown (6:00 am to 6:00 pm) will take place in the following areas:

  • SG-X4: Beamline Sectors: 12,13,14,15 and LOM 434
  • SG-X6: Beamline Sectors: 20,21,22,23 and LOM 436
  • SG-X7: Beamline Sectors: 24,25,26, and 27 (No LOM)

Please do not schedule any work in these areas that require electricity during the April/May shutdown until the final schedule has been set. More information will be coming during the run.

An email should have been received in regard to the X-Ray BPM Work during Dec/Jan Shutdown.

Sector 14: A new insertion device was added with hard-wired power limiting interlocks.
Concerns:
Sector 19: A device was removed/measured/reinstalled. No magnetic degradation was found.
Maintenance items:
Linear Encoders replaced on:

  • 21ds, 22, 23ds, 30us

Sector 34: Device removed/reinstalled to allow repair of a vacuum leak.

Users should check their beamline limits so that they can be raised or lowered if needed. Floor Coordinators may change these limits. Also, check for ID control access, particularly any new machines or user names.

If you make copies of your assigned ESAFS prior to posting, please make an extra copy for the Floor Coordinator at the same time. (i.e. if you give them two signed copies instead of one, it keeps them from doing unnecessary run around.)

If you need help, please call MOM. This site is available from the main AES page.

Gas Cylinder Storage, Kevin Beyer

Dr. Rosner has not returned to inspect the gas cylinder storage situation. However, he is still expected to do so. We have a total of 400 bottles stored. If you have cylinders that need to be returned, or if you need more storage, please contact Kevin Beyer. Kevin is in the process of implementing a system where an inventory of gas cylinders will be available through him to rectify the current storage problem.

Carl Nelson is working on repairs to the gas cylinder storage cage doors.

ESAF Upgrades January 2007, Bruce Glagola

Experimenter's Tab: Training information is displayed when the Find link is used.
Description Tab: New ESAF cannot have start date prior to current sysdate.
New Electrical Equipment Inspection Tab

Next Meeting:

The next CAT Chat meeting will be held on Monday, February 5, 2007.